# A DbC Inspired Neurosymbolic Layer for Trustworthy Agent Design

### v1

@@ 1 Introduction
Lastly, inspired by observational equivalence [26, 18, 31, 9], agents satisfying the same probabilistic contracts are functionally equivalent with respect to those contracts; they differ only in (i) $P_{succ}$, (ii) operational costs, and (iii) potential—i.e., the capacity of the agent to satisfy an ever more complex set of conditions, enabling principled comparison and selection.
@@ 3 Implementation
The contract layer is built entirely on the SymbolicAI framework [13] , extending its neurosymbolic capabilities with DbC-inspired validation mechanisms [^1] The contract source code is available at this GitHub [link]; the documentation is available at this GitBook [link].. At its core, the implementation leverages user-defined data models (Pydantic-based [7] LLMDataModel subclasses) that serve as the concrete instantiation of the type system $\mathcal{T}$. These models define not only structural requirements but also rich semantic descriptions through field annotations, guiding both type validation and LLM generation.
Figure 1: Contract Execution Flow
@@
The contract execution flow proceeds through distinct phases: (i) initial type validation of inputs against $\mathcal{T}$, (ii) pre-condition checking with optional remediation, (iii) an optional act method for intermediate transformations, (iv) LLM-driven output generation guided by the class’s prompt property and type specifications, and (v) post-condition validation with remediation attempts. When remediation is enabled for either pre- or post-conditions, both input and output objects are "fixed"—guaranteed to be type-valid with all field conditions satisfied. This fixing process uses the LLM to populate or correct fields according to the type schema and semantic instructions, ensuring that each field contains valid data (e.g., an extracted email or default value) that passes all specified constraints.
Remediation is achieved through ValidationFunction component that iteratively refines outputs by incorporating validation error messages into corrective prompts. The accumulation of error history across retry attempts enables the LLM to learn from previous failures, avoiding cyclic errors. Each validation function is parameterized by retry configurations in $\Theta$, controlling the maximum attempts, delays, and backoff strategies.
A critical aspect of the implementation is the fallback mechanism, which embodies DbC’s invariant principle. The original forward method is always executed in a finally block, regardless of contract validation outcomes. This ensures system resilience: when contract validation fails, the forward method can implement graceful falling by returning a safe, type-compliant default. This design guarantees that contract failures never prevent system operation, merely degrading from verified to best-effort behavior.
Lastly, the success of contract satisfaction for each invocation is a Bernoulli random variable, with the empirical success probability $P_{succ}$ computed over a sliding window of invocations. Combined with performance metrics tracking validation overhead and remediation costs, this enables runtime comparison of functionally equivalent agents that satisfy the same contracts $\mathcal{C}$.
@@ 4 Limitations and Future Work
Formal Verification. Current contracts lack formal guarantees about type system correctness. Future work will pursue Lean4 [27] formalization of the entire pipeline, providing machine-checked proofs of type safety and contract satisfaction properties. This formalization will establish rigorous foundations for trustworthy agent design, enabling verification of contract consistency, type safety preservation across agent compositions, and probabilistic bounds on contract satisfaction under various operational conditions.

### v2

@@ 1 Introduction
Inspired by observational equivalence [32, 22, 38, 13], agents satisfying the same probabilistic contracts are functionally equivalent with respect to those contracts; they differ only in (i) $P_{succ}$, (ii) operational costs, and (iii) potential—i.e., the capacity of the agent to satisfy an ever more complex set of conditions, enabling principled comparison and selection.
In practice, guardrails and schema validators [35, 11], function-calling and constitutional-style self-remediation [27, 9, 5, 28], and ReAct/RAG correction loops [49, 3, 43, 48, 7] can all be expressed as pre/post predicates with bounded remediation under the contract layer. Our contribution is a unified, declarative DbC formalism and execution flow that subsumes these patterns while remaining implementation-agnostic.
@@ 3 Implementation
The contract layer is built entirely on the SymbolicAI framework [17] , extending its neurosymbolic capabilities with DbC-inspired validation mechanisms [^1] The contract source code is available at this GitHub [link]; the documentation is available at this GitBook [link]. For challenging practical use cases involving contracts, see this codebase [link], which includes various types of contracts ranging from simple to complex (e.g., repairing a broken ontology).. At its core, the implementation leverages user-defined data models (Pydantic-based [11] LLMDataModel subclasses) that serve as the concrete instantiation of the type system $\mathcal{T}$. These models define not only structural requirements but also rich semantic descriptions through field annotations, guiding both type validation and LLM generation.
Figure 1: Contract execution flow.
@@
The contract execution flow proceeds through distinct phases: (i) initial type validation of inputs against $\mathcal{T}$, (ii) pre-condition checking with optional remediation, (iii) an optional act method for intermediate transformations, (iv) LLM-driven output generation guided by the class’s prompt property and type specifications, and (v) post-condition validation with remediation attempts. Type validation enforces structural well-formedness, while pre/post enforce semantic predicates on well-typed objects. When remediation is enabled for either pre- or post-conditions, both input and output objects are "fixed"—guaranteed to be type-valid with all field conditions satisfied. This fixing process uses the LLM to populate or correct fields according to the type schema and semantic instructions, ensuring that each field contains valid data (e.g., an extracted email or default value) that passes all specified constraints.
Remediation is achieved through TypeValidationFunction component that iteratively refines outputs by incorporating validation error messages into corrective prompts. The accumulation of error history across retry attempts enables the LLM to learn from previous failures, avoiding cyclic errors. Each validation function is parameterized by retry configurations in $\Theta$, controlling the maximum attempts (to preclude deep loops), delays, and backoff strategies.
A critical aspect of the implementation is the fallback mechanism, which embodies DbC’s invariant principle. The original forward method is always executed in a finally block, regardless of contract validation outcomes. If contract validation fails, the forward method can enable graceful falling, where contract or type mismatches do not raise—the raw forward output is returned as-is, or can return a safe, type-compliant default. This design guarantees that contract failures never prevent system operation, merely degrading from verified to best-effort behavior.
Lastly, the success of contract satisfaction for each invocation is a Bernoulli random variable. We compute the empirical success probability over N independent runs as $P_{\text{succ}}=\tfrac{1}{N}\sum_{t=1}^{N}\mathbf{1}[\text{all contract predicates pass}]$. To simplify evaluation and reporting, we approximate the overall success probability by multiplying per-family success probabilities. For example, if a post-condition involves six individual checks (predicates), we can group them into families—such as class existence, index validity, and cluster reduction—and treat these families as independent, even though dependencies may exist. Combined with performance metrics that track validation overhead and remediation costs, this approximation enables runtime comparison of functionally equivalent agents that satisfy the same contracts $\mathcal{C}$.
@@ 4 Limitations and Future Work
Formal Verification. Current contracts lack formal guarantees about type system correctness. Future work will pursue Lean4 [33] formalization of the entire pipeline, providing machine-checked proofs of type safety and contract satisfaction properties. This formalization will establish rigorous foundations for trustworthy agent design, enabling verification of contract consistency, type safety preservation across agent compositions, and probabilistic bounds on contract satisfaction under various operational conditions. The Lean4 effort is prospective; the runtime layer does not depend on it.